HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

FIRST ANALYSIS OF THE 0 − ν′′ PROGRESSION OF THE COMET-TAIL SYSTEM A^{2}Π_{i} − X^{2}Σ^{+} IN THE^{12}C^{17}O^{+} CATION.

I. PIOTROWSKA, W. SZAJNA, S. RYZNER, A. STASIK, A. PARA, P. KOLEK, R. KȨPA, \underline{R. HAKALLA} , Materials Spectroscopy Laboratory, Institute of Physics, University of Rzeszów, Pigonia 1 Street, 35-310 Rzeszów, Poland; M.I. MALICKA, The Faculty of Mathematics and Applied Physics,

Rzeszów University of Technology, Powstańców Warszawy 8 Street, 35-959 Rzeszów, Poland

The 0 − ν′′ progression of the Comet-Tail A^{2}Π_{i} − X^{2}Σ^{+} system, where ν′′ = (0, 1, 2, 3), was recorded for the first time in the ^{12}C^{17}O^{+} isotopologue. Using a high-resolution Fourier-transform spectrometer, more than 1500 emission lines were measured in the region of 12,880 - 20,500 cm^{−1}.

A deperturbation analysis was performed using the PGOPHER program.^{1} As a result, deperturbed molecular parameters of the A^{2}Π_{i}(v = 0) and X(v = 0,1,2,3) levels and the A^{2}Π_{i}(v = 0) ∼ X^{2}Σ^{+}(v = 10) perturbation parameter of a complex (spin-orbit, spin-electronic and rotation-electronic) character were determined.

The values of the ro-vibronic terms of the A^{2}Π_{i}(v = 0) level were also calculated and the percentage^{2}Π character of the A^{2}Π_{i}(v = 0) and X^{2}Σ^{+}(v = 10) levels was examined. This work is a continuation of the studies on the A^{2}Π_{i} state in the^{12}C^{17}O^{+} isotopologue made by our team.^{2}

1doi:10.1016/j.jqsrt.2016.04.010, C. M. Western, PGOPHER: a program for simulating rotational, vibrational and electronic spectra, Journal of Quantitative Spectroscopy and Radiative Transfer, 186, 221-242, (2017).

2doi:10.1016/j.jqsrt.2022.108268, I. Piotrowska, R. Hakalla, W. Szajna, R. Kȩpa, First observation of the Comet-Tail A^{2}Π_{i} − X^{2}Σ^{+} system in the^{12}C^{17}O^{+} cation, Journal of Quantitative Spectroscopy and Radiative Transfer, 289, 108268, (2022).

p-number: p121  Submitted on Wed, 14 Jun 2023 10:20:04 +0200